OpenStudy (anonymous):

Write [logbase3(7)][logbase4(3)][logbase7(6)] as a single logarithm.

OpenStudy (anonymous):

Write \[ (\log _{3}7)(\log _{4}3)(\log _{7}6) \] as a single logarithm